# To Trust or Not to Trust: On Calibration in ML-based Resource Allocation for Wireless Networks

### v1

@@ V-A Results
Figs. 6 and 7 presents the OP for a 4-resource system using OLF and BCE, respectively, across a range of $q_{th}$ values, for (a) $\gamma_{th}=0.5$ at $\mathtt{SNR}=2~{}\mathrm{dB}$ and (b) $\gamma_{th}=1.0$ at $\mathtt{SNR}~{}=~{}8~{}\mathrm{dB}$. Unlike Fig. 5 from [46], these plots use $q_{th}$ values on a logarithmic scale and also present the OP obtained using the calibrated predictors. It can be seen that OP obtained using OLF and BCE exhibit a ‘U’ shaped curve, indicating that both very small and very large $q_{th}$ values are not good. This is because very small $q_{th}$ values results in the user almost always being allocated the final resource, while very large $q_{th}$ values lead to allocation to the first resource. Notably, at $q_{th}=0.5$, the uncalibrated OLF-trained predictor outperforms the BCE-trained ones. As $q_{th}$ decreases, this performance gap narrows, and in some cases, BCE slightly outperforms OLF at lower $q_{th}$.
Similar to the observation made in Fig. 4, it can also be seen from Fig. 6 that the calibrated and uncalibrated OLF-trained predictors both achieve the same minimum OP but at different $q_{th}$ values. For instance, in a 4-resource system with $\gamma_{th}=0.5$ and $\mathtt{SNR}=2~{}\mathrm{dB}$ (see Fig. 6 (a)), the uncalibrated predictor trained with OLF achieves its minimum OP at $q_{th}=0.5$, which matches its training threshold, while the calibrated predictor reaches the same minimum OP at $q_{th}=10^{-3}$. This observation aligns with the result of Theorem 2, which establishes that post-processing calibration cannot reduce the minimum achievable OP. It is important to note that this observation is not reflected in Fig.7, since $q_{th}$ is not a hyperparameter in BCE-based training. As a result, Theorem 2 is not applicable here.
@@ V-B Proof of Concept
To demonstrate the application of the resource allocation technique and the OLF discussed in Section II[71] showcases video transmission over fading wireless channels using four different outage predictors. The goal is to compare the effectiveness of these predictors in reducing the likelihood of video outages and ensuring smooth playback. In this setup, a resource refers to a wireless channel used for video transmission. When a channel’s signal quality degrades due to fading and can no longer support reliable video transmission, an outage is likely to occur. However, the system leverages the outage predictor to anticipate such events and proactively switch to a more reliable channel. Playback is interrupted only when a predictor incorrectly recommends a channel that fails to meet the transmission requirement, in which case the screen turns red to indicate the outage (see Fig 8 (a)).
@@
This demonstration reinforces the effectiveness of the OLF-trained predictor, which resulted in an average outage rate of only 0.1% over the entire video duration, significantly lower than that of the the dummy model (6.8%), and the predictors trained using BCE (3.2%) and MSE (3.1%), leading to smoother video playback. Please note that the demonstration here uses the uncalibrated predictor with $q_{th}$ set to 0.5. However, it is worth highlighting that the calibrated predictor, trained using the OLF, will also achieve at best the same OP performance through the selection of an appropriate $q_{th}$, in accordance with Theorem 2 presented in this work. This behavior is consistent with Fig. 6, which shows that both the calibrated and uncalibrated predictors trained with OLF achieve the same minimum system OP, but at different $q_{th}$ (classification threshold) values.
@@ A-A Proof of Theorem 1
    $P_{\infty}(\gamma_{th},q_{th})$ $=\mathbb{P}\left[C<\gamma_{th}\mid Q^{\star}\leq q_{th}\right]$ (37)
@@
    For an $|\mathcal{R}|$ resource system: Substituting (40) and (36) in  [46, eq. (13)], we obtain (18). This completes the proof.
@@ A-C Proof of Theorem 3
$\noindent\hskip 1.00006pt\mathbb{E}[A(Q^{\star})\mid Q^{\star}\leq q_{th}]$ $<P$
@@
$\hskip 1.00006pt~{}~{}~{}+\mathbb{E}[A(Q^{\star})]\left(1-(1-F(q_{th}))^{|\mathcal{R}|-1}\right)$

### v3

@@ V-B Proof of Concept
To demonstrate the application of the resource allocation technique and the OLF discussed in Section II[73] showcases video transmission over fading wireless channels using four different outage predictors. The goal is to compare the effectiveness of these predictors in reducing the likelihood of video outages and ensuring smooth playback. In this setup, a resource refers to a wireless channel used for video transmission. When a channel’s signal quality degrades due to fading and can no longer support reliable video transmission, an outage is likely to occur. However, the system leverages the outage predictor to anticipate such events and proactively switch to a more reliable channel. Playback is interrupted only when a predictor incorrectly recommends a channel that fails to meet the transmission requirement, in which case the screen turns red to indicate the outage (see Fig. 10 (a)).
@@
This demonstration reinforces the effectiveness of the OLF-trained predictor, which resulted in an average outage rate of only 0.1% over the entire video duration, significantly lower than that of the dummy model (6.8%), and the predictors trained using BCE (3.2%) and MSE (3.1%), leading to smoother video playback. Please note that the demonstration here uses the uncalibrated predictor with $q_{th}$ set to 0.5. However, it is worth highlighting that the calibrated predictor, trained using the OLF, will also achieve at best the same OP performance through the selection of an appropriate $q_{th}$, in accordance with Theorem 2 presented in this work. This behavior is consistent with Fig. 6, which shows that both the calibrated and uncalibrated predictors trained with OLF achieve the same minimum system OP, but at different $q_{th}$ (classification threshold) values.
@@ VI Conclusion
Several fronts of new investigations can be envisaged for future work. One of them shall focus on exploring conformal prediction as an alternative to post-processing calibration methods. Unlike methods that recalibrate predicted confidence scores to better align with observed outcome frequencies, conformal prediction provides formal calibration guarantees, ensuring that the true outcome falls within a set of possible outcomes with a user-specified confidence level. This makes it particularly appealing in scenarios involving distribution shift. Another one shall explore systems operating in fading scenarios other than the Rayleigh one to understand how worse or better propagation conditions impact the calibration performance. A preliminary investigation appears in [2]. On another front, the aim is to extend the single-user approach used here to a multiuser-multirate approach, which better conforms to more practical and advanced networks.
@@ A-A Proof of Theorem 1
    $P^{\star}_{\infty}(\gamma_{th},q_{th})$ $=\mathbb{P}\left[C<\gamma_{th}\mid Q^{\star}\leq q_{th}\right]$ (38)
@@
    For an $|\mathcal{R}|$ resource system: Substituting (41) and (37) in  [48, eq. (13)], we obtain (19). This completes the proof.
@@ A-C Proof of Theorem 3
$\noindent\thinspace\mathbb{E}[A(Q^{\star})\mid Q^{\star}\leq q_{th}]$ $<P$
@@
$\thinspace~~~+\mathbb{E}[A(Q^{\star})]\left(1-(1-F(q_{th}))^{|\mathcal{R}|-1}\right)$